A football is kicked at a velocity of 15 m/s at an angle of 25 o to the horizontal. Calculate:

a) The total flight time

b) The maximum height

c) The range of the ball.


1
Expert's answer
2020-10-19T13:20:55-0400

a) Let's write the equation of motion of the ball in horizontal and vertical directions:


"x=v_0tcos\theta, (1)""y=v_0tsin\theta-\dfrac{1}{2}gt^2, (2)"

here, "x" is the horizontal displacement of the ball (or the range of the ball), "v_0 = 15 \ \dfrac{m}{s}" is the initial velocity of the ball, "t" is the total flight time of the ball, "\theta=25^{\circ}" is the launch angle, "y" is the vertical displacement of the ball (or the height) and "g = 9.8 \ \dfrac{m}{s^2}" is the acceleration due to gravity.

We can find the total flight time of the ball from the second equation (since ball returns to the ground, "y = 0"):


"0=v_0tsin\theta-\dfrac{1}{2}gt^2,""t = \dfrac{2v_0sin\theta}{g}=\dfrac{2\cdot 15\ \dfrac{m}{s}\cdot sin25^{\circ}}{9.8 \ \dfrac{m}{s^2}}=1.29 \ s."

b) Let's first find the time that ball takes to reach the maximum height from the kinematic equation:


"v_y = v_0sin\theta - gt_{rise},""0=v_0sin\theta - gt_{rise},""t_{rise} = \dfrac{v_0sin\theta}{g}."

Then, we can substitute "t_{rise}"into the second equation and find the maximum height:


"y_{max}=v_0sin\theta\cdot\dfrac{v_0sin\theta}{g}-\dfrac{1}{2}g(\dfrac{v_0sin\theta}{g})^2,""y_{max}=\dfrac{v_0^2sin^2\theta}{2g} = \dfrac{(15\ \dfrac{m}{s})^2 \cdot sin^2 25^{\circ}}{2\cdot 9.8 \ \dfrac{m}{s^2}} = 2.05 \ m."

c) Finally we can substitute the total flight time into the first equation and find the range of the ball:


"x=v_0cos\theta \dfrac{2v_0sin\theta}{g} = \dfrac{v_0^2sin2\theta}{g},""x=\dfrac{(15\ \dfrac{m}{s})^2 \cdot sin2\cdot 25^{\circ}}{9.8 \ \dfrac{m}{s^2}}=17.59\ m."

Answer:

a) "t =1.29 \ s."

b) "y_{max}= 2.05 \ m."

c) "x=17.59\ m."